`If <1 and <2 are complementary angles and m<1= (2x+20) and m<2= (3x +15) find the value of x?
Answer by stanbon(75887)

`If <1 and <2 are complementary angles and m<1= (2x+20) and m<2= (3x +15) find the value of x?
----
Equation::
(2x+20) + (3x+15) = 90 degrees
------
5x + 35 = 90
5x = 55
x = 11
----------------------
<1 = 2x+20 = 22+20 = 42 degrees
<2 = 3x+15 = 33+15 = 48 degrees
-----------
Cheers,
Stan H.
-------------